OpenStudy (lena772):

Relative Maximum at ? if f'(x)=2(x-1)((2x+1)^2)(5x-2)

OpenStudy (lena772):

x = 1 only x = 2/5 only x = 1/7 only x = -1/2 only x = 1 and x = -1/2

hero (hero):

You could find f''(x) to figure this out, but it appears that it might be more of a hassle than just multiplying everything out then integrating f'(x) to find f(x). Once you find f(x) you could simply graph it and find the relative max.